OpenStudy (anonymous):

the capacity of a cuboidal tank is 50000 liters of water.find the breadth of the tank,if its length and depth are respectively 2.5m and 10m.

OpenStudy (anonymous):

2m is d answer

OpenStudy (bahrom7893):

50 000 liters = 50 meters cubed 50 = l*b*h 50 = 2.5*10*b 50 = 25b b = 2m

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@Parv: Give Bahrom a medal.

OpenStudy (anonymous):

k

OpenStudy (bahrom7893):

lol tnx veng and parv